Feb 3, 2021 Calorimeter
Meeting Day/Time: every other WEDNESDAY at 11:00 a.m. JLab time
- To join via a Web Browser, go to the page  https://bluejeans.com/907185247.
- To join via Polycom room system go to the IP Address: 22.214.171.124 (bjn.vc) and enter the meeting ID: 907185247.
- To join via phone, use one of the following numbers and the Conference ID: 907185247.
- US or Canada: +1 408 740 7256 or
- US or Canada: +1 888 240 2560
- Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
- More information on connecting to bluejeans is available.
Participant Direct Lines
- JLab Phone in CC L207 is 757-269-7084 (usual room)
- JLab Phone: in CC F326 is 757-269-6460
- Phone in the Regina Video-conference Suite is 306-585-4204
- Athens Phone: in Christina's office is 011-30-210-727-6947
- FCAL HDFCAL log book
- BCAL HDBCAL log book
- Calibrations: 2020 Data Production; RunPeriod-2019-11 Validation; Offline Monitoring Data Validation
Goals for Calorimetry Group
- Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
- Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at least 5 MeV.
- Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.
- Run Preparations
- Z calibration (K Suresh) BCAL Z Update
- Peaking of Noise Showers at 220 cms in the BCAL
- For both the Photon Gun sample and the bggen sample, Run number 50715 was used. One sample was generated without any BKG (BKG=None) and another with (BKG=Random:recon-2018_08-ver02). Version XML : recon-2018_08-ver02_9.xml
- Tolga: Slight update to Photon Reconstruction checks for BCAL/Question
- Action Items
- Any other business
Attending: Tolga, Karthik, Sean, Zisis, Mark, Rebecca, Varun, Jon, Churamani, Matt, Elton, Susan
- The CALWG meeting scheduled for Feb 17 will be cancelled, as the Feb Collaboration meeting starts that afternoon and has a heavy schedule, including calorimeter updates from PrimEx. The next CALWG meeting will thus be on March 3, perhaps jointly with PrimEx.
- Task Force reports at Collab Meeting: Sean will give a summary and technical reports will be written. One more photon recon meeting will be scheduled for next week.
- Matt mentioned that our collaboration has a draft charge for the March Review.
- Zisis reported that the COMCAL paper is being edited by the co-authors.
- Run Preparations
- FCAL: Mark will check on when the next 100 bases will come out for refurbishing. Firmware testing is ongoing. Hovanes and Nick will work on the integration within CSS. Jon et al will be consulted on the remote uploading procedure to the bases. (At press: The first 100 bases are back in the FCAL now. )
- BCAL: no report.
- Karthik reported on z-position correction to the non-linearity of the BCAL z-calibration. Currently, we only have E-cor implemented, so he is picking up z-cor from this studies last year. As expected, for high Egamma the z-coverage in the BCAL gets squeezed to large z values. Uses a single z bin to look at a correction. Karthik generated his own bggen sample instead of using the large-statistics one available for 2 reasons: a) the latter contains random noise (see issue in next bullet point below) and b) it does not contain the uncorrected/raw energies in the trees, which he needs for this study. He used the pi0 skims and compared to Mark's gamma-gun MC. Sean recommended that the fits and background modelling be improved possibly by increasing the window around the pi0.
- Karthik shows a spike of random noise showers at 212 cm in the BCAL. Mark suggested that this could be due to how timing is assigned to these randoms and Sean agreed. Although the spike would not affect Karthik's studies (it can be cut out) it could be an issue for cross section extraction and it needs to be investigated further. Karthik will enter it as a GitHub issue so that Richard can have a look.
- BCAL LED
- Varun showed an update on the SiPM dark rate study from the 2020 run period. The plots were shown layer wise; channel-wise plots will be made next to see which ones are driving an increase of the pedestal RMS at the higher biases.
- BCAL (offline): Tolga had no update; he is waiting for batch 10.
- FCAL (offline): (Susan) 2019-11 monitoring of FCAL: Batch 9 looks ok (only one extra DeltaT band). Susan showed her analysis on the FCAL LED, called 'LED Monitoring' (vs the old 'bad channels' name). She proposed a new way of doing the efficiency ('present approach') which results in efficiencies =<1; the group accepted this. In the 'regions' plots 0 is not pedestal but events with gain all the way down. LED 'regions' will be renamed 'LED ranges' and reflect that the LED color and bias settings cause the sum over the ADC integrals to end up in specific histogram ranges. Matt/Mark suggested imposing a minimum integral cut. 2017 LED data is different and efficiencies already exist so is low priority (scripts will need adjustment); 2017 checks were done and things are good. Mark will wait for Spring 2018 numbers before updating the db. Rebecca could use those new numbers to generate MC, although the differences should be small. Next Susan will look at LED triggers versus data triggers.
- BCAL LED
- Efficiency: Jon can a brief update on using the b1 omega-pi for BCAL efficiency studies; Karthik will check further on its feasibility (statistics, and 3pi inv mass background).
- Reconstruction: Tolga gave update to Photon Reconstruction checks for BCAL/Question. A discussion on the U1 flag ensued and Sean explained its use: without it trees only have particles that make combos. Tolga will continue checks.
- Simulations: no report.
- Action Items: BCAL z=212 cm issue added to Short Term items and Karthik will add it to GitHub.
- Any other business: none.